# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

### v1

@@ V-C Generative AI and Distributed Learning
In FedGen [181], the server based on local model performance cooperatively trains a GAI model as a lightweight generator to amass user information and that is capable of generating pseudo data representation without accessing their true data (see Fig. 8). Thus, each local user can have the additional knowledge of other distributed clients, without sharing data among distributed devices. This approach is robust as this innovative approach can significantly enhance the performance and communication efficiency of FL, all while ensuring that user data remain secure and private. In [44], an AaaS framework based on AIGC models is designed to address the challenges in training and deployment overheads for edge networks.
Similar to FedGen, which leverages GAI to train a generative AI model, DENSE [182] designs a GAN-based data generation network that enables the generation of on-server data. In contrast to FedGen, where GAI is employed for on-server data generation, DENSE employs a multi-teacher knowledge distillation technique (see Fig. 9). This technique ensure that the generated data accurately captures the entire system’s data distribution and characteristics. The local models collaboratively instruct a student, denoted as the global model, inheriting knowledge from all teachers. Furthermore, this approach can be enhanced by incorporating Label-Driven Knowledge Distillation (LKD) [183]. This methodology enables the student to selectively learn the most meaningful knowledge from each teacher, significantly improving distillation efficiency compared to other conventional knowledge distillation approaches. Consequently, it is feasible to apply the ensemble learning on the server, boosting the FL efficiency while retaining the communication efficiency.
Different from other GAN approaches, the work in [184] adopts a distinct approach by treating model parameters as data. Specifically, the authors introduce HCFL, a technique capable of compressing and reconstructing model parameters. It enables the FL system to circumvent substantial communication overheads. In particular, applying HCFL to FL systems can maintain stable accuracy when the distortion rate of the aggregated model decreases proportionally with network size.
@@
Aside from the aforementioned methods, invariant learning is a unique approach to enhancing learning on local devices in distributed learning. In invariant learning, users collect their local invariant representations or cross-domain invariant representations by sharing the most consistent format of data (invariant features) across devices. In this way, it minimizes the need for extensive communication resources, a representative form of knowledge-aided SemCom in distributed learning. For instance, Federated Cross-Correlation and Continual Learning (FCCL) [185] and FCCL+ [186] are innovative methods inspired by Barlow Twins principles, which help identify invariants across domains in the FL system (See Fig. 10). As an alternative method to discover invariant features, the authors in [187] focus on class-wise invariance by introducing cluster prototypes of contrastive learning. Unlike [185, 186, 187], FedPAC employs a structured causal model to acquire goal-oriented invariant representations, demonstrating robustness in learning personalized representations for each client without extensive knowledge sharing among devices [188].
@@
Table V: GAI for Optimizing Semantic Communication Scenarios.
@@
Apps Ref Model Inputs Outputs Loss function Performance
@@
Knowledge Abstraction [141] Transformer-based GAN Textual input source Summarized knowledge Cross Entropy Competitive results on document retrieval with less memory
[136] BERT GAN Language Data Summarized Knowledge Log-likelihood Only necessitates a basic definition of relations as input, making it appropriate for extracting knowledge about novel, intricate relations that are immediately established and were previously unavailable
@@
[189] LERP Knowledge Graph with missing edges Contextual information as vectors Gradient-based Outperforms other rule learning and estimates expressive logical rules
[190] ADT-GAN Multivariate time series data Anomaly scores for each data point Adversarial Loss Effectiveness in anomaly detection within multivariate time series data
@@
 [22] LLMs Data and information from edge networks Collaborative planning and decision-making by multi-agent LLMs MSE Demonstrates the potential benefits of on-device LLMs to solve network intents.
@@
 [191] DGMs Distorted reconstructions of images Improved perceptual quality of transmitted images MSE Improves perceptual quality by performing denoising using style-based GAN
Data Compression  [192] VQ-VAE Spectral envelope parameters from speech signals Improved quantized spectral envelope parameters for speech vocoder Vector quantization Reduces log spectral distortion and increases perceptual evaluation of speech quality
@@
 [193] AE-LSTM Heterogeneous data from IoT devices Identification of out-of-norm activities for cyber threat hunting in IIoT MSE Ensemble model outperforms conventional ML with better compression
@@
 [94] ELAN Captured images from edge devices Semantic optimized resource allocation extracted from images MSE Reduces transmission costs by extracting important semantic information
@@
 [148] NECST Heterogeneous data from IoT devices Improved quality reconstructed images MSE Significant gain in data reconstruction under different settings of channel distortion
@@
 [157] DeepSC Captured images from edge devices Realistic image reconstructed based on semantic segmentation BLEU Immense resource gain in bandwidth reduction during semantic segment transmission
@@
 [167] RC Captured images from edge devices Lossless reconstructed images. Cross-entropy Lossless image reconstructed from lossy compressed data
@@
 [179] Causal semantic communication Sequential signal Reconstructed sequential signal KL Divergence Improves signal reconstruction using causal theorem to predict the potential signal to support the reconstruction
@@
 [177, 178] NeSy Graph representation Knowledge based on graph KL Divergence Using neural-symbolism to improve the AI task of reconstructing knowledge from the graph
@@
Semantic Resource Allocation  [194] DRL Real-time data from vehicular networks Improved service quality, reliable guidance to receiving vehicles MSE Enhances service quality, reliability, and efficiency in GAI-enabled vehicular networks
@@
 [195] ISGC User inputs for semantics and graphics rendering Optimized resource allocation and high-quality content generation Variational autoencoder Optimal resource allocation strategy for the Metaverse
@@
 [196] Diffusion Wireless network collaboration tasks Optimized execution of AIGC tasks MSE Optimized edge computation resource utilization
@@
 [197] RSA Multi-modal data from human/machine agents Locally rendered scenes and interactions in the Metaverse MSE Utilizes distributed learning and multi-agent RL for efficient communication
@@
Distributed Learning  [22] LLMs Edge networks and multi-agent system parameters Collaboratively planned and solved tasks in wireless networks Custom ensemble loss Enables self-governed networks with on-device LLMs for intelligent decision-making
@@
 [43] FL-GAN AIGC tasks, settings of wireless networks, and privacy Diverse, personalized, and high-quality content generation Adversarial loss Improved learning efficiency, privacy protection, and content quality
@@
[181] FL-GAN Heterogeneous user data Ensembled knowledge for guiding local models KL Divergence Improved generalization performance in FL with minimum communication resource
@@
 [198] Diffusion Network scale, environment parameters, and user demands Customized network solutions and expert-free optimization Custom ensemble loss Best adaptation to objectives and scenarios and creation of novel designs
V-D Lessons Learned and Keys Takeaway
@@
In summary, the key takeaways from this section are as follows.
  • GAI markedly enhances compression performance through the creation of concise and organized data representations.
  • GAI empowers the SemCom capability to abstract knowledge effectively, facilitating high-performance data compression without sacrificing information.
  • GAI facilitates the generation of pseudo data on distributed devices, allowing these devices to adapt to the domains of other devices in distributed learning scenarios.
  • GAI can assist the encoder in generating invariant features from data, comprising the most meaningful features that contribute significantly to AI performance. This improvement leads to enhanced learning while concurrently achieving a substantial reduction in communication costs.
  • While current knowledge abstraction heavily relies on NLP models, there exists significant research potential in SemCom. There is a pressing need for the introduction of a generalized, scalable approach to abstract knowledge across types, eliminating the necessity of intermediate processes like NLP.
@@ VI Challenges and Future Directions
GAI has demonstrated its effective abilities in the development of mobile networks, such as enhancing the user experience, personalizing services, optimizing network performance, and enabling new business models. In return, it also poses several new issues that need to be addressed. In what follows, we will explore and propose numerous viable solutions to technical challenges of GAI (Section VI-A), including lack of interpretability (Section VI-B); ethical and legal concerns (Section VI-C); and security, privacy, and trust (Section VI-D). Fig. 11 illustrates the categories of open research challenges and directions prevalent in mobile networking through GAI. Finally, we summarize all critical issues and the respective solutions in Table VI.
Figure 11: Open research directions in mobile networking through GAI.

### v2

@@ V-C Generative AI in Distributed Learning
Similar to FedGen, which leverages GAI to train a GAI model, DENSE [203] designs a GAN-based data generation network that enables the generation of on-server data. In contrast to FedGen, where GAI is employed for on-server data generation, DENSE employs a multi-teacher knowledge distillation technique (see Fig. 10). This technique ensure that the generated data accurately captures the entire system’s data distribution and characteristics. The local models collaboratively instruct a student, denoted as the global model, inheriting knowledge from all teachers. Furthermore, this approach can be enhanced by incorporating Label-Driven Knowledge Distillation (LKD) [204]. This methodology enables the student to selectively learn the most meaningful knowledge from each teacher, significantly improving distillation efficiency compared to other conventional knowledge distillation approaches. Consequently, it is feasible to apply the ensemble learning on the server, boosting the FL efficiency while retaining the communication efficiency.
Unlike other GAN approaches, the work in [205] adopts a distinct approach by treating model parameters as data. Specifically, the authors introduce HCFL, a technique capable of compressing and reconstructing model parameters. It enables the FL system to circumvent substantial communication overheads. In particular, applying HCFL to FL systems can maintain stable accuracy when the distortion rate of the aggregated model decreases proportionally with network size.
@@
Aside from the aforementioned methods, invariant learning is a unique approach to enhancing learning on local devices in distributed learning. In invariant learning, users collect their local invariant representations or cross-domain invariant representations by sharing the most consistent format of data (invariant features) across devices. This way, it minimizes the need for extensive communication resources, a representative form of knowledge-aided SemCom in distributed learning. For instance, Federated Cross-Correlation and Continual Learning (FCCL) [206] and FCCL+ [207] are innovative methods inspired by Barlow Twins principles, which help identify invariants across domains in the FL system (See Fig. 11). As an alternative method to discover invariant features, the authors in [208] focus on class-wise invariance by introducing cluster prototypes of contrastive learning. Unlike [206, 207, 208], FedPAC employs a structured causal model to acquire goal-oriented invariant representations, demonstrating robustness in learning personalized representations for each client without extensive knowledge sharing among devices [209].
@@
Table VI: GAI for Optimizing Semantic Communication Scenarios.
@@
Apps Ref Model Inputs Outputs Loss function Performance Limitations
@@
Knowledge Abstraction [154] Encoder-Decoder Textual input source Summarized knowledge Softmax Competitive results on document retrieval with less memory Required pre-processing of training data, and only applicable to text-based data applications
@@
[210] Logical Network Knowledge Graph with missing edges Contextual information as vectors Gradient-based Outperforms other rule learning and estimates expressive logical rules Hard to deploy and fine-tune LERP for various networking uses
[187] VAE Video and motion sensor measurements Localization Data VIB + APP Effectiveness in localization task-oriented data Caused straggling due to channel feedback requirements
@@
 [24] LLMs Data and information from edge networks Collaborative planning and decision-making by multi-agent LLMs Q-reward loss Demonstrates the potential benefits of on-device LLMs to solve network intents. The impractical algorithm due to the overwhelming computation overheads
@@
 [211] StyleGAN Distorted reconstructions of images Improved perceptual quality of transmitted images MSE + GAN loss Improves perceptual quality by performing denoising using style-based GAN High training complexity, mode collapse, and slow convergence rate
Data Compression  [180] Vision Transformer Image input source Improved quantized vector with noise mask Cross-entropy Generate simulated semantic noise to combat channel noise via adversarial training Not suitable for IoT frameworks due to the modeling over-parameterizations
@@
 [212] AE-LSTM Heterogeneous data from IoT devices Identification of out-of-norm activities for cyber threat hunting in IIoT MSE Ensemble model outperforms conventional ML with better compression Not applicable to different types of datasets
@@
 [106] Transformer and Diffusion Captured images from edge devices Semantic optimized resource allocation extracted from images MSE Reduces transmission costs by extracting important semantic information Not suitable for IoT systems due to the overparameterized network
@@
 [131] Diffusion model Heterogeneous data from IoT devices Improved quality reconstructed images Log-likelihood Significant gain in data reconstruction under different settings of channel distortion Required large test times and computation overheads due to the intensive diffusion process
@@
 [169] Transformer Captured images from edge devices Realistic image reconstructed based on semantic segmentation BLEU Immense resource gain in bandwidth reduction during semantic segment transmission Not feasible for IoT systems due to transformer architecture, high computation, and large model size
@@
 [213] RC Captured images from edge devices Lossless reconstructed images. Cross-entropy Lossless image reconstructed from lossy compressed data Unable to simulate information loss of LBP at different scales
@@
 [200] Causal semantic communication Sequential signal Reconstructed sequential signal KL Divergence Improves signal reconstruction using causal theorem to predict the potential signal to support the reconstruction Information loss due to neglected non-causal components
@@
 [198, 199] NeSy Graph representation Knowledge based on graph KL Divergence Neural-symbolism to improve the AI task of reconstructing knowledge from the graph Straggling due to the requirement to receive feedback from the receiver to improve the data compression
@@
Semantic Resource Allocation  [214] DRL Real-time data from vehicular networks Improved service quality, reliable guidance to receiving vehicles MSE Enhances service quality, reliability, and efficiency in GAI-enabled vehicular networks Required large test times and computation overheads due to the intensive diffusion process
@@
 [215] ISGC User inputs for semantics and graphics rendering Optimized resource allocation and high-quality content generation Variational autoencoder Optimal resource allocation strategy for the Metaverse Required large test times and computation overheads due to the intensive diffusion process
@@
 [216] Diffusion model Wireless network collaboration tasks Optimized execution of AIGC tasks MSE Optimized edge computational resource utilization Required large test times and computation overheads due to the intensive diffusion process
@@
 [217] RSA Multi-modal data from human/machine agents Locally rendered scenes and interactions in the Metaverse MSE Utilizes distributed learning and multi-agent RL for efficient communication High communication overheads due to frequently shared semantic representations among agents
@@
Distributed Learning  [24] LLMs Edge networks and multi-agent system parameters Collaboratively planned and solved tasks in wireless networks Q-reward loss Self-governed networks with on-device LLMs for intelligent decision-making The impractical algorithm due to overwhelming computation overheads
@@
 [40] Diffusion models + LLM AIGC tasks, settings of wireless networks, and privacy Diverse, personalized, and high-quality content generation Adversarial loss Improved learning efficiency, privacy protection, and content quality Not suitable for IoT systems due to fine-tuning foundation models
@@
[202] GAN Heterogeneous user data Ensembled knowledge for guiding local models KL Divergence Improved generalization performance in FL with minimum communication resource Poor generalization in learning due to estimating the models’ representations from noise (at the server)
@@
 [218] Diffusion model Network scale, environment parameters, and user demands Customized network solutions and expert-free optimization Custom ensemble loss Best adaptation to objectives and scenarios and creation of novel designs Required large test times and computation overheads due to the intensive diffusion process
V-D Lessons Learned and Key Takeaways
@@
The key takeaways from this section are as follows.
  • Key 1. GAI markedly enhances compression performance via the creation of concise and organized data representations.
  • Key 2. GAI empowers the SemCom capability to abstract knowledge effectively, facilitating high-performance data compression without sacrificing information.
  • Key 3. GAI facilitates the generation of pseudo data on distributed devices, allowing these devices to adapt to the domains of other devices in distributed learning scenarios.
  • Key 4. GAI can assist the encoder in generating invariant features from data that contribute significantly to AI performance. This improvement leads to enhanced learning while concurrently achieving a substantial reduction in communication costs.
@@ VI Challenges and Future Directions
GAI has demonstrated its effective abilities in network management, wireless security, and SemCom for mobile and wireless networking developments. However, integrating GAI into modern networks also raises new issues that must be addressed. In what follows, we have explored some viable solutions for the issues of complexity of GAI models in Section VI-A; scalability in Section VI-B; security, privacy, and trust in Section VI-C; and standardization in Section VI-D. Finally, we summarize all critical issues, the respective solutions, and the raised research directions in Table VII.